# Re-parameterized Low-rank Prompt: Generalize a Vision-Language Model within 0.5K Parameters

### v2

@@ 5.1.1 Base-to-new Generalization
RLP 0.5K 91 images/s 738 images/s
5.1.2 Domain Generalization
Then we follow CoCoOp [55] to use ImageNet, ImageNet-A, ImageNet-R, ImageNet-v2, and ImageNet-S to run domain generalization experiments to verify the robustness of RLP. Results are shown in Tab. 3. As seen from the table, on target datasets, RLP gets comparable or even better average accuracy compared with the latest methods. Importantly, RLP uses much fewer parameters and thus can reach significantly better efficiency. For CLIP, CoOp, CLIP-Adapter and ProGrad, there is a clear performance gap between our RLP and them.
5.1.3 Cross-dataset Transfer
Finally, we follow CoCoOp [55] to conduct cross-dataset transfer evaluation. Results are shown in Tab. 4. We can see that on the source dataset, CoOp accomplishes the highest score, which is similar with the situation in the base-to-new generalization setting. However, concentrating too much on the current dataset will absolutely cause overfitting and catastrophic forgetting problems, and finally lead to a severe drop in the performance on those unseen datasets. In this setting, RLP wins on 6 of 10 datasets and its average accuracy is also slightly better than the best competitor CoCoOp. Such result well demonstrates that RLP could maximally extract general and data-agnostic knowledge from given images compared with other prompt-based methods. Considering the huge difference in the parameter numbers, we could summarize that RLP is still the better choice.
5.1.4 Few-shot Learning
In this paragraph, we will show the experiment results of RLP in the few-shot learning setting. This setting is originated from CoOp. Seen from Fig. 4, RLP consistently outperforms zero-shot CLIP, CoOp, and CLIP-Adapter across all the shot numbers. Such results demonstrate the superiority of RLP in adaptation ability when there are few samples in downstream tasks.
Overall, in base-to-new generalization, domain generalization, cross-dataset transfer and few-shot learning, the proposed method can consistently accomplish state-of-the-art performance while enjoying extremely high parameter efficiency, fruitfully demonstrating the effectiveness and efficiency of the proposed method.
5.2 Analysis
5.2.1 Ablation study
We first investigate the impact of each component in the proposed method RLP. We revise the baseline model, i.e., CoOp [54] which adopts full-rank tunable text prompt, step-by-step.
Our modifications start with transforming the text encoder. We first replace the tunable prompt tokens with our low-rank prompt tokens, denoted as ``Low-rank prompts'', for the CoOp method. After replacing the classic prompts with our low-rank prompts, the average Harmonic accuracy over 11 datasets directly improved from 71.66% to 75.46%. The accuracy on the base classes decreases by 2.78% and on the new classes increases by 8.26%. Although the adaptation ability for base classes slightly drops, the generalization ability for new classes raises quite a lot. This phenomenon proves that our low-rank design is fairly beneficial for the model's generalization ability once again.
Then we integrate the re-parameterized full-rank initialization into CoOp. The harmonic mean accuracy improves by 1.27%. Here we can see the importance of special initialization, even if we have already changed the prompt format from full-rank to low-rank.
Finally, we add a delicate selected lightweight regularization layer, Dropout, to the current module. Dropout helps us alleviate the overfitting problem, and avoid catastrophic forgetting. Therefore, we get a very competitive result, 76.53% Harmonic average accuracy, which is completely comparable to some latest lightweight state-of-the-art methods.
5.2.2 Efficiency Comparison
@@
In this paragraph, we will give a comprehensive analysis of all the training, storage, and inference efficiencies for RLP and several existing methods. Since the parameter scale of CLIP-Adapter [10] is significantly larger than others, we do not contain CLIP-Adapter into comparison.
@@
For a fair comparison, we do all the speed tests on the same GPU. Results are shown in Tab. 6. Our proposed RLP shares nearly the same fastest training and inference speeds with the simplest method CoOp, and more importantly, RLP merely uses 0.5K parameters, which is a lot more storage-efficient than other methods. Specially, compared with classic method CoCoOp, RLP enjoys $>$18x training speed, $>$56x inference speed and $<$70x storage usage. Besides complex structures, the huge gap in the inference speed is partly owing to the huge memory cost of CoCoOp, which forces us to adopt a smaller batch size than other methods for CoCoOp. Moreover, compared with the latest method ProGrad, RLP enjoys $>$1.6x training speed, comparable inference speed, and $<$60x storage usage, which adequately demonstrates the super efficiency of RLP.
Figure 5: Top line: Results of different training epochs on base-to-new generalization. Bottom line: Results of different dropout ratios on base-to-new generalization.
Table 7: Average results on 11 datasets for CLIP with convolutional image encoder, ResNet-50 [11].
image encoder #params base new H
CoOp ResNet-50 2.1K 77.16 61.01 68.14
ProGrad ResNet-50 8.2K 73.29 65.96 69.06
RLP ResNet-50 0.5K 75.22 64.98 69.73
Table 8: Results of adding RLP to image prompts on ImageNet.
#params base new H
Text RLP 0.5K 75.87 70.67 73.17
Image RLP 0.5K 74.57 69.40 71.89
5.2.3 Results for Convolutional Image Encoder
In this subsection, we will show the experiment results of RLP on CLIP with convolutional image encoder ResNet-50 [11], rather than the default ViT-B/16 [7]. Seen from Tab. 7, compared with baseline CoOp, RLP still largely improves the new accuracy and the harmonic mean accuracy over 11 datasets, while the base accuracy slightly drops. Compared with the latest method ProGrad, RLP shows clear superiority on base accuracy and the harmonic mean accuracy.
5.2.4 Results for Using RLP on the Image Side
There are several works [18, 21] indicating the effectiveness of the image prompts. Therefore, in this subsection, we will explore the results of applying RLP to image prompts. Seen from Tab. 8, using RLP on the image side also reaches high accuracy. However, the base, new, and average accuracy of image RLP is not as good as those of text RLP. Such experiment result tells us to use RLP on the text side instead of the image side when we aim to reach high accuracy with extremely few parameters.
5.2.5 Results across Different Hyper-parameters
Effect of Different Epochs
In this paragraph, we investigate that how the total training epoch could influence the adaptation result. As shown in Fig. 5, we select a series of epoch numbers, i.e. 25, 50, 75, 100, and 200. And we run an experiment using RLP to get the corresponding results for the given epoch. We could see that as the training epoch increases, the accuracy on base classes continues decreasing while the accuracy on new classes continues increasing. It is reasonable because as the training continues, the model has a higher and higher risk of forgetting its original knowledge and overfitting the current dataset. To reach a better harmonic mean accuracy, a good trade-off is necessary.
Effect of Different Dropout Ratios
In this paragraph, we will show the influence of different dropout ratios in RLP on ImageNet. Seen from Fig. 5, as the dropout ratio increases, the base accuracy starts decreasing while the new accuracy starts increasing in most cases. The harmonic mean accuracy first increases and then decreases. Such a result is reasonable. Since a high dropout ratio can be regarded as a kind of strong regularization, proper regularization will help avoid overfitting and catastrophic forgetting problems. However, too strong regularization will damage the ability of the model.
Table 9: Results of different ranks on ImageNet.
rank #params base new H
1 0.5K 75.87 70.67 73.17
2 1.0K 76.03 70.13 72.96
3 1.5K 76.20 70.40 73.19
lr wd Acc 1e-4 5e-4 1e-3
1e-3 70.73 70.67 70.78
2e-3 70.80 70.83 70.77
3e-3 70.83 70.83 70.81
Effect of Different Ranks
In this paragraph, we will show the influence of different ranks in RLP. Seen from Tab. 9, roughly, a larger rank brings more parameters, higher base accuracy, and lower new accuracy. As a result, increasing rank is not necessarily able to improve the average accuracy.
Effect of Different Learning Rates and Weight Decays
In this paragraph, we investigate the effect of normal hyper-parameters learning rate and weight decay. We wonder if tuning them carefully would lead to significant improvement. Seen from Tab. 10, the performance of RLP stays stable whatever the learning rate and weight decay vary. The conclusion here is that our method RLP is robust for learning rate and weight decay.
6 Conclusion
With the development of huge vision-language models, how to effectively and efficiently adapt such huge models to various downstream tasks becomes a challenging problem. Much effort has been made to leverage the potential of prompt tuning in adapting vision-language models. However, existing methods usually suffer from inefficiency. To solve such a problem, we propose Re-parameterized Low-rank Prompts (RLP) for a better trade-off between effectiveness and efficiency in vision-language adaptation. Along with the low-rank prompt design, we further propose a novel initialization method and integrate the Dropout layer as a lightweight regularization module to further improve the performance without introducing any extra parameters and inference cost. We are the first to explore how to effectively adapt vision-language models using an extremely small number of parameters, i.e. 0.5K. Besides efficiency and effectiveness, RLP has many other advantages such as simplicity and robustness, which are quite valuable in real-world application scenarios. Extensive experiments show that even if RLP uses much fewer parameters, RLP could still consistently be comparable to or better than state-of-the-art methods.

### v3

@@ 5.2 Analysis
5.2.1 Ablation study
We first show the impact of components of DIP step by step.
We start from transforming the text encoder. We first replace the tunable prompt tokens with our decomposed small prompt matrices, denoted as “Decomposition”, for the CoOp method. After such replacement, the average Harmonic accuracy over 11 datasets directly improved from 71.66% to 75.46%. The accuracy on the base classes decreases by 2.78% and on the new classes increases by 8.26%. Although the adaptation ability slightly drops, the generalization ability raises quite a lot. This phenomenon proves that our low-rank design is fairly beneficial for the model’s generalization ability once again. Then we integrate the special full-rank initialization into CoOp. The harmonic mean accuracy improves by 1.27%. Finally, we add a lightweight regularization layer, Dropout. It helps us alleviate overfitting and catastrophic forgetting, resulting in further improvement.
5.2.2 Additional experimental results
Effect of different training epochs In this paragraph, we investigate that how the total training epoch could influence the adaptation result. Shown in Fig. 4, we run experiments for the given epochs separately. As the training epoch increases, the accuracy on base classes continues decreasing while the accuracy on new classes continues increasing. It is reasonable because as the training continues, the model has a higher risk of forgetting its original knowledge and overfitting.
Effect of different dropout ratios In this paragraph, we will show the influence of different dropout ratios in DIP on ImageNet. Seen from Fig. 4, as the dropout ratio increases, the base accuracy starts decreasing while the new accuracy starts increasing mostly. The harmonic mean first increases and then decreases. Dropout iss a kind of regularization, only proper regularization can help avoid overfitting and catastrophic forgetting.
CLIP with convolutional image encoder In this paragraph, we show the results of DIP on CLIP with convolutional image encoder ResNet-50 [57], rather than the default ViT-B/16 [51]. Seen from Section 5.2.2, compared with baseline CoOp, DIP still largely improves the new accuracy and the harmonic mean accuracy over 11 datasets, while the base accuracy slightly drops. Compared with the latest method ProGrad, DIP shows clear superiority on base accuracy and the harmonic mean accuracy.
Table 6: Results on ResNet-50 encoded CLIP.
#params base new H
CoOp 2.1K 77.16 61.01 68.14
ProGrad 8.2K 73.29 65.96 69.06
DIP+CoOp 0.5K 75.22 64.98 69.73
Table 7: Effect of DIP’s rank on ImageNet.
rank #params base new H
1 0.5K 75.87 70.67 73.17
2 1.0K 76.03 70.13 72.96
3 1.5K 76.20 70.40 73.19
Table 8: Results of adding DIP to image prompts on ImageNet.
#params base new H
Text DIP 0.5K 75.87 70.67 73.17
Image DIP 0.5K 74.57 69.40 71.89
Effect of different ranks In this paragraph, we will show the influence of different ranks in DIP. Seen from Section 5.2.2, roughly, a larger rank brings more parameters, higher base accuracy, and lower new accuracy. As a result, increasing rank is not necessarily able to improve the average accuracy.
Results for using DIP on the image prompts There are several works [25, 28] indicating the effectiveness of the image prompts. Therefore, in this subsection, we will explore the results of applying DIP to image prompts. Seen from Tab. 8, using DIP on the image side also reaches high accuracy. However, the base, new, and average accuracy of image DIP is not as good as those of text DIP. Such experiment result tells us to use DIP on the text side instead of the image side when we aim to reach high accuracy with extremely few parameters.
6 Conclusion
With the development of huge vision-language models, how to effectively and efficiently adapt such huge models to downstream tasks becomes a challenging problem. Much effort has been made to leverage the potential of prompt tuning in adapting vision-language models. However, existing methods suffer from inefficiency. To reach extremely efficient generalization, we propose Dense Information Prompt (DIP) based on the observation about the strong correlation between prompt rank and generalizability. Moreover, we propose a novel initialization method and a lightweight regularization module to further improve the low-rank design without adding any extra inference cost. We are the first to explore how to effectively adapt vision-language models using an extremely small number of parameters, i.e. 0.5K. Besides efficiency and effectiveness, DIP has many valuable advantages such as simplicity and robustness. Extensive experiments and analyses show the superiority of DIP sufficiently.
Appendix A Datasets
Following previous work [7, 6], we leverage 11 image recognition datasets to verify the effectiveness of the proposed method for both the base-to-new generalization task. These datasets include two datasets for the generic object classification, i.e., ImageNet [58] and Caltech101 [59], five datasets for the fine-grained classification, i.e., OxfordPets [60], StanfordCars [61], Flowers102 [62], Food101 [63] and FGVCAircraft [64], one dataset for the scene recognition, i.e., SUN397 [53], one dataset for the action recognition, i.e., UCF101 [65], one dataset for the texture classification, i.e., DTD [54], and one dataset for the satellite imagery recognition, i.e., EuroSAT [66]. Following previous works [6, 7], for each dataset, we split its classes equally into two non-overlapping groups, i.e., one as base classes and the other as new classes. We train all models on the base classes and perform a base/new evaluation on the base/new classes.
Appendix B Training Details
To verify the effectiveness of our proposed method, we explore the improvement of integrating DIP into a lightweight prompt tuning method CoOp and a heavy prompt tuning method MaPLe separately.
For DIP+CoOp and DIP+MaPLe, we conduct a grid search to find the optimal hyper-parameters based on the configuration of CoOp and MaPLe. In the main text, we set the rank $r=1$ in DIP for all the experiments unless specially mentioned. The total number of tunable parameters in DIP+CoOp is $4*1+1*512=516$, which is merely around 0.25x of that in CoOp. For DIP+MaPLe, we also decompose its weights of the projection layer that projects text prompts to generate image prompts with rank $r_{proj}=64$. Therefore, the total number of tunable parameters in DIP+MaPLe is $(4*1+1*512)+(512*64+64*768)=82436$ per layer, and $82436*9=741924$ in total for 9 layers are modified in MaPLe by default.
Table 9: Detailed comparisons with latest methods in base-to-new generalization. H: harmonic mean [55]. DIP+CoOp outperforms all the other lightweight methods on 8 out of 11 datasets, and DIP+MaPLe achieves better harmonic mean accuracy on 9 out of 11 datasets compared to MaPLe alone. Better still, methods modified by DIP even enjoys significantly better parameter efficiency.
((a)) Parameters comparison..
Method CoOp CoCoOp Adapter LoRA ProGrad DIP+CoOp MaPLe DIP+MaPLe
#params 2.1K 35.4K 525.5K 129.0K 8.2K 0.5K 3548K 741.9K
((b)) Average.
Base New H
CLIP 69.34 74.22 71.70
CoOp 82.69 63.22 71.66
CoCoOp 80.47 71.69 75.83
Adapter 82.62 70.97 76.35
LoRA 84.30 67.33 74.86
ProGrad 82.79 68.55 75.00
DIP+CoOp 80.32 74.73 77.42
MaPLe 82.28 75.14 78.55
DIP+MaPLe 83.17 75.43 79.11
((c)) ImageNet.
Base New H
CLIP 72.43 68.14 70.22
CoOp 76.47 67.88 71.92
CoCoOp 75.98 70.43 73.10
Adapter 76.53 66.67 71.26
LoRA 74.77 58.47 65.62
ProGrad 77.03 68.80 72.68
DIP+CoOp 76.10 71.17 73.55
MaPLe 76.66 70.54 73.47
DIP+MaPLe 76.80 70.97 73.77
((d)) Caltech101.
Base New H
CLIP 96.84 94.00 95.40
CoOp 98.00 89.81 93.73
CoCoOp 97.96 93.81 95.84
Adapter 98.20 93.20 95.63
LoRA 98.49 90.33 94.24
ProGrad 98.50 91.90 95.09
DIP+CoOp 98.10 95.23 96.65
MaPLe 97.74 94.36 96.02
DIP+MaPLe 98.20 94.83 96.49
((e)) OxfordPets.
Base New H
CLIP 91.17 97.26 94.12
CoOp 93.67 95.29 94.47
CoCoOp 95.20 97.69 96.43
Adapter 94.40 94.10 94.25
LoRA 94.90 92.57 93.72
ProGrad 94.40 95.10 94.75
DIP+CoOp 95.53 97.93 96.72
MaPLe 95.43 97.76 96.58
DIP+MaPLe 95.67 98.00 96.82
((f)) StanfordCars.
Base New H
CLIP 63.37 74.89 68.65
CoOp 78.12 60.40 68.13
CoCoOp 70.49 73.59 72.01
Adapter 77.13 69.23 72.97
LoRA 81.07 65.30 72.34
ProGrad 79.00 67.93 73.05
DIP+CoOp 71.33 74.47 72.87
MaPLe 72.94 74.00 73.47
DIP+MaPLe 76.00 73.00 74.47
((g)) Flowers102.
Base New H
CLIP 72.08 77.80 74.83
CoOp 97.60 59.67 74.06
CoCoOp 94.87 71.75 81.71
Adapter 97.70 70.83 82.13
LoRA 98.23 60.20 74.65
ProGrad 96.27 71.07 81.77
DIP+CoOp 95.53 75.23 84.18
MaPLe 95.92 72.46 82.56
DIP+MaPLe 97.00 73.77 83.80
((h)) Food101.
Base New H
CLIP 90.10 91.22 90.66
CoOp 88.33 82.26 85.19
CoCoOp 90.70 91.29 90.99
Adapter 90.40 90.40 90.40
LoRA 88.57 87.30 87.93
ProGrad 90.17 89.53 89.85
DIP+CoOp 90.83 92.03 91.43
MaPLe 90.71 92.05 91.38
DIP+MaPLe 90.63 91.90 91.26
((i)) FGVCAircraft.
Base New H
CLIP 27.19 36.29 31.09
CoOp 40.44 22.30 28.75
CoCoOp 33.41 23.71 27.74
Adapter 39.57 32.27 35.55
LoRA 46.27 28.83 35.53
ProGrad 42.63 26.97 33.04
DIP+CoOp 35.17 35.57 35.37
MaPLe 37.44 35.61 36.50
DIP+MaPLe 39.03 37.17 38.08
((j)) SUN397.
Base New H
CLIP 69.36 75.35 72.23
CoOp 80.60 65.89 72.51
CoCoOp 79.74 76.86 78.27
Adapter 81.67 73.93 77.61
LoRA 79.73 69.00 73.98
ProGrad 80.70 71.03 75.56
DIP+CoOp 79.60 77.80 78.69
MaPLe 80.82 78.70 79.75
DIP+MaPLe 81.27 78.13 79.67
((k)) DTD.
Base New H
CLIP 53.24 59.90 56.37
CoOp 79.44 41.18 54.24
CoCoOp 77.01 56.00 64.85
Adapter 80.47 52.23 63.35
LoRA 82.93 54.90 66.06
ProGrad 76.70 46.67 58.03
DIP+CoOp 75.67 60.43 67.20
MaPLe 80.36 59.18 68.16
DIP+MaPLe 82.10 59.50 69.00
((l)) EuroSAT.
Base New H
CLIP 56.48 64.05 60.03
CoOp 92.19 54.74 68.69
CoCoOp 87.49 60.04 71.21
Adapter 86.93 64.20 73.86
LoRA 94.90 65.67 77.62
ProGrad 91.37 56.53 69.85
DIP+CoOp 82.70 65.40 73.04
MaPLe 94.07 73.23 82.35
DIP+MaPLe 93.67 74.33 82.89
((m)) UCF101.
Base New H
CLIP 70.53 77.50 73.85
CoOp 84.69 56.05 67.46
CoCoOp 82.33 73.45 77.64
Adapter 85.80 73.63 79.25
LoRA 87.47 68.03 76.53
ProGrad 83.90 68.50 75.42
DIP+CoOp 82.93 76.70 79.69
MaPLe 83.00 78.66 80.77
DIP+MaPLe 84.50 78.13 81.19
Appendix C Competitors
  1. 1.
    CLIP [1]: CLIP is a strong baseline vision-language model that is pre-trained on a large number of image-text pairs from the web by learning a contrastive objective. CLIP enables strong zero-shot adaptation ability on various downstream tasks by using fixed text prompts, i.e. a photo of a.
  2. 2.
    CoOp [7]: CoOp replaces the fixed text prompts in CLIP with tunable text prompts to improve the adaptation ability of the vision-language model. CoOp shows excellent performance in few-shot situations.
  3. 3.
    CoCoOp [6]: CoCoOp replaces the isolated tunable text prompts in CoOp with conditional text prompts, which receive extra gradients from the image features besides text features. CoCoOp largely improves the generalization ability of the vision-language model, getting good results on base-to-new generalization and domain adaptation.
  4. 4.
    CLIP-Adapter [56]: CLIP-Adapter adopts the thoughts of classic Adapter [36] to use serial linear layers and activation functions to adapt for downstream tasks. It is simple yet effective in few-shot learning.
  5. 5.
    LoRA [37]: LoRA adopts low-rank decomposition for weights in FC layers. It is efficient and earns good results in NLP field.
  6. 6.
    MaPLe [28]: MaPLe simultaneously adds prompts to the image encoder and text encoder of CLIP. To trigger more information exchange between the image side and text side, MaPLe generate image prompts from the projection of text prompts. Though effective, such design brings quite heavy cost.
  7. 7.
    ProGrad [52]: ProGrad only updates the text and image prompts whose gradient are aligned (or non-conflicting) to the general knowledge, which is represented as the optimization direction offered by the pre-defined prompt predictions. Such regularization helps it finish good adaptation and generalization.
Table 10: Base-to-new generaliation performances based on SLIP [71].
#params base new H
CoOp 2.1K 68.45 42.77 52.64
DIP+CoOp 0.5K 62.53 47.78 54.17
lr wd Acc 1e-4 5e-4 1e-3
1e-3 70.73 70.67 70.78
2e-3 70.80 70.83 70.77
3e-3 70.83 70.83 70.81
Table 12: Deep prompts in different depths for DIP+CoOp.
depth #params base new H
1 0.5K 76.37 74.69 75.52
2 1.0K 77.81 73.87 75.79
3 1.5K 79.55 72.72 75.98
4 2.0K 80.07 72.69 76.20
5 2.6K 80.24 73.37 76.65
6 3.1K 80.64 73.39 76.85
Appendix D Complete Results in Base-To-New Generalization
As shown in Tab. 9, overall, DIP shows quite competitive performance, and the superiority mainly relies on the improvement of new class accuracy. In other words, DIP largely improves the generalization ability of CLIP. In particular, compared with CoOp, DIP+CoOp gets $10.95$% accuracy gain on the new classes and $2.83$% accuracy drop on the base classes, showing strongest harmonic mean accuracy on 8 out of 11 datasets among all the lightweight existing methods. Compared with a heavy method MaPLe, DIP+MaPLe achieves better harmonic mean accuracy on 9 out of 11 datasets.
Appendix E Experiments on a Different Vision-Language Architecture
In this paragraph, we show the results of DIP on another vision-language architecture, SLIP [71], besides CLIP. Seen from Tab. 10, DIP+CoOp earns much higher new accuracy and harmonic mean accuracy than the original CoOp.
Appendix F Effect of Deep Prompts
In this paragraph, we extend the shallow prompts in DIP+CoOp to the deep prompts. We record the accuracy change as we increase the layers including tunable prompts, following the last equation in Section 3.1 in the main text. Results are shown in Tab. 12. As the depth increases, the base accuracy keeps increasing while the new accuracy first decreases and then increases. Overall, more depth generally leads to higher harmonic mean accuracy. Therefore, it is possible to further improve the performance of our method by increasing the prompt depth.
Appendix G Effect of Different learning rates and Weight Decays
In this paragraph, we investigate the effect of normal hyper-parameters learning rate and weight decay. We wonder if tuning them carefully would lead to significant improvement. Seen from Tab. 11, the performance of DIP stays stable whatever the learning rate and weight decay vary. The conclusion here is that our method DIP is robust for learning rate and weight decay.
@@ Appendix G Effect of Different learning rates and Weight Decays
DIP+CoOp 0.5K 91 images/s 738 images/s
Appendix H Efficiency Comparison
@@
In this paragraph, we will give a comprehensive analysis of all the training, storage, and inference efficiencies for DIP and several existing methods. Since the parameter scale of CLIP-Adapter [56] is significantly larger than others, we do not contain CLIP-Adapter into comparison.
@@
For a fair comparison, we do all the speed tests on the same GPU. Results are shown in Tab. 13. Our proposed DIP shares nearly the same fastest training and inference speeds with the simplest method CoOp, and more importantly, DIP merely uses 0.5K parameters, which is a lot more storage-efficient than other methods. Specially, compared with classic method CoCoOp, DIP enjoys $>$18x training speed, $>$56x inference speed and $<$70x storage usage. Besides complex structures, the huge gap in the inference speed is partly owing to the huge memory cost of CoCoOp, which forces us to adopt a smaller batch size than other methods for CoCoOp. Moreover, compared with the latest method ProGrad, DIP enjoys $>$1.6x training speed, comparable inference speed, and $<$60x storage usage, which adequately demonstrates the super efficiency of DIP.